# Unbounded-width CSPs are Untestable in a Sublinear Number of Queries

### v1

@@ C.2 Concentration of Value for Random Instances
Recall from Definition 4.3 that, in process of generating the instance $I\sim\mathcal{D}_{\mathrm{no}}^{\mathsf{G},n,d}$, the total of $dn$ constraints are placed in $d$ rounds. In each round, for each $j\in[n]$, exactly one constraint involving $(j,1)$ is added. This naturally divides the stochastic process of generating $I$ into $dn$ time steps, where exactly one constraint is added to $I$ in each step. Let $\mathscr{F}_{t}$ be the $\sigma$-algebra corresponding to the information after $t$ time steps are completed, for each $t\in\{0,1,\dots,dn\}$. For each fixed assignemnt $\tau:[n]\times[3]\rightarrow\mathsf{G}$, we define the random variable $X_{\tau}:=\mathrm{val}_{I}(\tau)$ and apply Proposition C.3 to $X_{\tau}$ and the filtration $(\mathscr{F}_{t})_{0\leqslant t\leqslant dn}$. Using a “switching” argument (e.g., see [Wor99, Lemma 2.19]), it is not hard to see that
@@
for any $t\in[dn]$. Furthermore, we have $\mathbb{E}\left[X_{\tau}|\mathscr{F}_{dn}\right]=X_{\tau}=\mathrm{val}_{I}(\tau)$ and $\mathbb{E}\left[X|\mathscr{F}_{0}\right]=\mathbb{E}\left[X_{\tau}\right]=|\mathsf{G}|^{-1}$. Therefore, the conclusion of Proposition C.3 implies
@@
$\underset{i\in[n]}{\mathbb{P}}\left[\tau\big((i,x)\big)=\pi(x)\right]\geqslant\textup{Perm-val}(\tau)$ (C.2)
@@
We first show the existence of a map $\pi:D\to D$ satisfying (C.2) for all $x\in D$, and subsequently prove that the map $\pi$ thus obtained is a permutation.
@@
$\underset{(i,x)\in[n]\times D}{\mathbb{P}}\left[\tau\big((i,x)\big)\neq y_{x}\right]>2\varepsilon$ (C.3)
@@
Case 1: $\textnormal{Perm-val}(\tau)\geqslant 1-2\varepsilon$. Since $2\varepsilon\leqslant\frac{1}{5|D|}$, we may apply Lemma C.6 and obtain a permutation $\pi:D\rightarrow D$ such that
@@
holds for any $x\in D$. If $R\big((\pi(x))_{x\in D}\big)=1$, then this contradicts the assumption that (C.3) holds for all $y\in D^{D}$ with $R(y)=1$. Therefore, we must have $R\big((\pi(x))_{x\in D}\big)=0$.
@@
Case 2: $\textnormal{Perm-val}(\tau)<1-2\varepsilon$. Since $R(y)\leqslant\textit{Perm}_{D}(y)$ for all $y\in D^{D}$, by Definition C.5 we have
@@
Appendix D Varieties Admitting Affine-type
@@ D.1 Varieties
D.2 Types of Strictly Simple Algebras
Following conventions in universal algebra, we refer to an algebraic structure simply as an “algebra.” An algebra of $k$ elements is said to be simple if $k\geqslant 2$ and every homomorphic image of it has either 1 element or $k$ elements. It is further said to be strictly simple if every subalgebra of it has either 1 element or $k$ elements. Tame Congruence Theory, developed by Hobby and Mckenzie [HM88], classified strictly simple algebras into five type: (1) the unary type, (2) the affine type, (3) the Boolean type, (4) the lattice type and (5) the semilattice type. Later, Szendrei [Sze92] provided a characterization of strictly simple idempotent algebras of all five types. We record the following immediate corollary of [Sze92, Theorem 6.1]; readers interested in further details are referred to [HM88, Sze92].
Proposition D.3 ([Sze92]).
@@ D.2 Types of Strictly Simple Algebras
In general, a finite algebra (not necessarily strictly simple) is associated with a subset of the five possible types {unary, affine, Boolean, lattice, semilattice}. A variety $\mathcal{V}$ is said to admit a type if that type appears in the type set of at least one algebra belonging to $\mathcal{V}$. The following theorem by Valeriote [Val09] is a crucial component of Theorem 3.2.
Theorem D.5 ([Val09, Proposition 3.1]).
@@
The final ingredient in Theorem 3.2 is the following characterization of bounded-width CSP templates in terms of the variety generated by its polymorphism algebra. It was conjectured by [LZ07] and proved by [BK14].
Theorem D.6 ([BK14, Conjecture 4.3]).
@@
It is then clear that Theorem 3.2 follows from the combination of Corollary D.4 and Theorems D.5 and D.6.

### v2

@@ D.2 Concentration of Value for Random Instances
Recall from Definition 4.3 that, in process of generating the instance $I\sim\mathcal{D}_{\mathrm{no}}^{\mathsf{G},n,d}$, the total of $dn$ constraints are placed in $d$ rounds. In each round, for each $j\in[n]$, exactly one constraint involving $(j,1)$ is added. This naturally divides the stochastic process of generating $I$ into $dn$ time steps, where exactly one constraint is added to $I$ in each step. Let $\mathscr{F}_{t}$ be the $\sigma$-algebra corresponding to the information after $t$ time steps are completed, for each $t\in\{0,1,\dots,dn\}$. For each fixed assignemnt $\tau:[n]\times[3]\rightarrow\mathsf{G}$, we define the random variable $X_{\tau}:=\mathrm{val}_{I}(\tau)$ and apply Proposition D.3 to $X_{\tau}$ and the filtration $(\mathscr{F}_{t})_{0\leqslant t\leqslant dn}$. Using a “switching” argument (e.g., see [wormald1999models, Lemma 2.19]), it is not hard to see that
@@
for any $t\in[dn]$. Furthermore, we have $\mathbb{E}\left[X_{\tau}|\mathscr{F}_{dn}\right]=X_{\tau}=\mathrm{val}_{I}(\tau)$ and $\mathbb{E}\left[X|\mathscr{F}_{0}\right]=\mathbb{E}\left[X_{\tau}\right]=|\mathsf{G}|^{-1}$. Therefore, the conclusion of Proposition D.3 implies
@@
$\underset{i\in[n]}{\mathbb{P}}\left[\tau\big((i,x)\big)=\pi(x)\right]\geqslant\textup{Perm-val}(\tau)$ (D.2)
@@
We first show the existence of a map $\pi:D\to D$ satisfying (D.2) for all $x\in D$, and subsequently prove that the map $\pi$ thus obtained is a permutation.
@@
$\underset{(i,x)\in[n]\times D}{\mathbb{P}}\left[\tau\big((i,x)\big)\neq y_{x}\right]>2\varepsilon$ (D.3)
@@
Case 1: $\textnormal{Perm-val}(\tau)\geqslant 1-2\varepsilon$. Since $2\varepsilon\leqslant\frac{1}{5|D|}$, we may apply Lemma D.6 and obtain a permutation $\pi:D\rightarrow D$ such that
@@
holds for any $x\in D$. If $R\big((\pi(x))_{x\in D}\big)=1$, then this contradicts the assumption that (D.3) holds for all $y\in D^{D}$ with $R(y)=1$. Therefore, we must have $R\big((\pi(x))_{x\in D}\big)=0$.
@@
Case 2: $\textnormal{Perm-val}(\tau)<1-2\varepsilon$. Since $R(y)\leqslant\textit{Perm}_{D}(y)$ for all $y\in D^{D}$, by Definition D.5 we have
@@
Appendix E Varieties Admitting Affine-type
@@ E.1 Varieties
E.2 Types of Strictly Simple Algebras
Following conventions in universal algebra, we refer to an algebraic structure simply as an “algebra.” An algebra of $k$ elements is said to be simple if $k\geqslant 2$ and every homomorphic image of it has either 1 element or $k$ elements. It is further said to be strictly simple if every subalgebra of it has either 1 element or $k$ elements. Tame Congruence Theory, developed by Hobby and Mckenzie [hobby1988structure], classified strictly simple algebras into five type: (1) the unary type, (2) the affine type, (3) the Boolean type, (4) the lattice type and (5) the semilattice type. Later, Szendrei [szendrei1992survey] provided a characterization of strictly simple idempotent algebras of all five types. We record the following immediate corollary of [szendrei1992survey, Theorem 6.1]; readers interested in further details are referred to [hobby1988structure, szendrei1992survey].
Proposition E.3 ([szendrei1992survey]).
@@ E.2 Types of Strictly Simple Algebras
In general, a finite algebra (not necessarily strictly simple) is associated with a subset of the five possible types {unary, affine, Boolean, lattice, semilattice}. A variety $\mathcal{V}$ is said to admit a type if that type appears in the type set of at least one algebra belonging to $\mathcal{V}$. The following theorem by Valeriote [valeriote2009subalgebra] is a crucial component of Theorem 3.2.
Theorem E.5 ([valeriote2009subalgebra, Proposition 3.1]).
@@
The final ingredient in Theorem 3.2 is the following characterization of bounded-width CSP templates in terms of the variety generated by its polymorphism algebra. It was conjectured by [larose2007bounded] and proved by [barto2014constraint].
Theorem E.6 ([barto2014constraint, Conjecture 4.3]).
@@
It is then clear that Theorem 3.2 follows from the combination of Corollary E.4 and Theorems E.5 and E.6.